# fMoE: Fine-Grained Expert Offloading for Large Mixture-of-Experts Serving

### v1

@@ 6.6 Sensitivity Analysis
We analyze the sensitivity of three hyperparameters: prefetch distance of MoE models, the capacity of Expert Map Store, and inference batch size.
@@
Prefetch distance of MoE models. Figure 13 shows the TTFT and TPOT of fMoE when serving three MoE models with different prefetch distances. We have demonstrated that the expert hit rate decreases when gradually increasing the prefetch distance (Figure 4). When the prefetch distance is small (), fMoE cannot perfectly hide its system delay from the inference process, such as the map matching and expert prefetching, leading to the increase of inference latency. With larger prefetch distances (), fMoE has worse expert hit rates that also degrade the performance. Therefore, we set the prefetch distance to 3 for evaluating fMoE.
@@
Capacity of Expert Map Store. We measure the mean semantic and trajectory similarity scores searched in fMoE’s expert map matching for MoE model serving. Figure 14(a) presents the mean semantic and trajectory similarity scores of fMoE with different Expert Map Store capacity sizes. Both semantic and trajectory similarity scores improve as the store capacity increases. While the similarity scores exhibit a significant increase with capacities below 1K, further capacity expansion yields diminishing similarity gains. To minimize fMoE’s memory overhead, we set fMoE’s Expert Map Store capacity to 1K in evaluation.
@@
Inference batch size. We investigate the impact of inference batch size on fMoE and three baselines using Mixtral-87B with LMSYS-Chat-1M. Figure 14(b) presents the performance of fMoE, Mixtral-Offloading, ProMoE, and MoE-Infinity as the batch size increases from one to four. fMoE achieves the lowest TTFT and TPOT in most cases.
6.7 System Overheads
@@
Latency overheads of fMoE’s operations. Figure 15 shows the latency breakdown of one inference iteration in fMoE when serving the three MoE models. We report any operations of fMoE in §6.2 that may incur a significant latency delay, including context collection, map matching, expert on-demand loading, expert prefetching, and map update after the iteration completes. Qwen1.5-MoE has lower end-to-end iteration latency than Mixtral-87B and Phi-3.5-MoE because of significantly fewer parameters. Note that expert prefetching, map matching, and map update tasks are executed asynchronously, aside from the inference process. Hence, they do not contribute to the end-to-end iteration latency. Excluding three asynchronous tasks, the total delay incurred by other operations is consistently less than 30ms (5% of the iteration) across three MoE models, which is negligible compared to the inference latency.
@@
Memory overheads of fMoE’s Expert Map Store. Figure 16 shows the CPU memory footprint of fMoE’s Expert Map Store when varying the store capacity from 1K to 32K maps. The memory needed to store expert maps for Qwen1.5-MoE is more than Mixtral-87B and Phi-3.5-MoE because it has more experts per layer over the other two models, which increases the map shape. Even for the largest capacity (32K), the Expert Map Store requires less than 200MB of memory to store the maps, which is trivial since modern GPU servers usually have abundant CPU memory (e.g., p4d.24xlarge on AWS EC2 [5] has over 1100 GB of CPU memory). In the evaluation, fMoE’s map store capacity with 1K maps is sufficient for maintaining performance (§6.6), resulting in minimal memory overhead.
7 Related Work
@@
Lossless MoE serving. Recent studies on lossless MoE serving have been widely proposed. DeepSpeed inference [4] offload layer-wise parameters without expert awareness. Mixtral-Offloading [16] employs LRU expert caching and introduces speculative prediction to enable expert prefetching. MoE-Infinity [54] proposes the request-level expert activation matrix to guide offloading in coarse granularity. SwapMoE [42] maintains a set of critical experts in GPU memory and adjusts them based on workload changes to minimize offloading overhead. ProMoE [47] trains predictors per MoE layer to achieve high speculative prediction accuracy and low inference latency. Lina [28] moves infrequently used experts to host memory and focuses more on distributed MoE training. Liu et al. [31] serves MoE models on serverless computing by predicting expert patterns with black-box Bayesian techniques. Unlike existing coarse-grained offloading solutions, fMoE tracks fine-grained expert patterns from both trajectory and semantic aspects and outperforms SOTA baselines.
Lossy MoE serving. Except for lossless offloading, other works also propose lossy MoE serving. Expert pruning [50, 13] reduces memory usage by removing underutilized experts. Knowledge distillation [43, 33] produces compact sparse MoE models. ComPEFT [56] demonstrates expert compression without accuracy loss, while MC-SMoE [39] further decomposes merged experts into low-rank and structurally sparse alternatives. Hobbit [49] uses low precision to serve less-critical experts. However, lossy serving may impact the generation quality and is orthogonal to fMoE.
@@
MoE refactorization. Some works propose to redesign and refactor the current MoE architecture. Pre-gated MoE [20] utilizes preemptive expert selection to eliminate the sequential dependencies between expert selection and execution. SiDA-MoE [14] proposes a sparsity-inspired, data-aware inference system that decouples the expert routing from inference. READ-ME [7] refactors pre-trained dense LLMs into specialized MoE models. Unlike the above works, fMoE requires zero training to serve open-source MoE models.
8 Conclusion

### v2

@@ 6.7. Sensitivity Analysis
We analyze the sensitivity of prefetch distance of MoE models, Expert Map Store capacity, and inference batch size.
@@
Prefetch distance of MoE models. Figure 15 shows the TTFT and TPOT of FineMoE when serving three MoE models with different prefetch distances. We have demonstrated that the expert hit rate decreases when gradually increasing the prefetch distance (Figure 4). When the prefetch distance is small, FineMoE cannot perfectly hide its system delay from the inference process, such as the map searching and expert prefetching, leading to an increase in inference latency. With larger prefetch distances, FineMoE has worse expert hit rates that also degrade performance. Therefore, we set the prefetch distance to 3, 6, and 4 for Mixtral-87B, Qwen1.5-MoE, and Phi-3.5-MoE, respectively.
@@
Capacity of Expert Map Store. We measure the mean semantic and trajectory similarity scores searched in FineMoE’s expert map searching for MoE model serving. Figure 16(a) presents the mean semantic and trajectory similarity scores of FineMoE with different Expert Map Store capacity sizes. Both semantic and trajectory similarity scores improve as the store capacity increases. While the similarity scores exhibit a significant increase with capacities below 1K, further capacity expansion yields diminishing similarity gains. To minimize FineMoE’s memory overhead, we set FineMoE’s Expert Map Store capacity to 1K in evaluation.
@@
Inference batch size. We investigate the impact of inference batch size on FineMoE and three baselines using Mixtral-87B with LMSYS-Chat-1M. Figure 16(b) presents the performance of FineMoE, Mixtral-Offloading, ProMoE, and MoE-Infinity as the batch size increases from one to eight. FineMoE achieves the lowest TTFT and TPOT in most cases.
6.8. System Overheads
We measure and report the system overheads of FineMoE.
@@ 6.8. System Overheads
Latency overheads of FineMoE’s operations. Figure 17 shows the latency breakdown of one inference iteration in FineMoE when serving the three MoE models. We report operation overheads of FineMoE, including context collection, map searching, expert on-demand loading, expert prefetching, and map update after the iteration completes. Qwen1.5-MoE has lower end-to-end iteration latency than Mixtral-87B and Phi-3.5-MoE because of significantly fewer parameters. Note that expert prefetching, map searching, and map update tasks are executed asynchronously, aside from the inference process. Hence, they do not contribute to the end-to-end iteration latency. Excluding three asynchronous tasks, the total delay incurred by other operations is consistently less than 50ms (1% of the iteration) across three MoE models, which is negligible compared to the inference latency.
@@
Memory overheads of FineMoE’s Expert Map Store. Figure 18 shows the CPU memory footprint of FineMoE’s Expert Map Store when varying the store capacity from 1K to 32K maps. The memory needed to store expert maps for Qwen1.5-MoE is more than Mixtral-87B and Phi-3.5-MoE because it has more experts per layer over the other two models, which increases the map shape. Even for the largest capacity (32K), the Expert Map Store requires less than 200MB of memory to store the maps, which is trivial since modern GPU servers usually have abundant CPU memory (e.g., p4d.24xlarge on AWS EC2 (AWS, 2006) has over 1100 GB of CPU memory). In evaluation, FineMoE’s map store capacity with 1K maps is sufficient for maintaining performance (§6.7), resulting in minimal memory overhead.
7. Discussion
In this section, we compare the heuristic-based FineMoE with Neural Network (NN)-based predictors, analyze the impact of model parallelism on FineMoE’s performance, and discuss how FineMoE can be extended to other MoE architectures.
NN-based predictors. NN-based predictors for expert offloading are impractical due to multiple sources of overhead. First, they often introduce sub-second inference latency, comparable to MoE inference latency itself. Second, they require extensive data collection, hour-long per-layer training, and frequent retraining to adapt to workload shifts. Third, they consume substantial GPU memory, as prior work (Song et al., 2024) reports millions of parameters per MoE layer. Moreover, they are incompatible with FineMoE ’s fine-grained design: training on fine-grained data hinders convergence, while storing iteration-level probabilities generates large volumes that further prolong training and limit feasibility. Therefore, we adopt a heuristic-based design rather than NN-based approaches.
Impact of EP and tensor parallelism (TP). Higher EP distributes experts across more devices and enables greater expert replication, which can increase FineMoE ’s offloading opportunities and memory savings. In contrast, higher TP raises the overhead of offloading operations, since dense model components are split across devices and require coordinated offloading and reloading. As noted in prior work (Liu et al., 2024a; Dai et al., 2024a), production MoE systems generally avoid high TP because its communication costs outweigh performance benefits. In large-scale deployments (e.g., DeepSeek (Liu et al., 2024a)), MoE systems usually use low EP during prefill to maximize throughput, while adopting high EP during decode to enable higher expert redundancy. Though high-EP decode reduces per-GPU expert occupancy, the larger number of expert replicas (e.g., 2 more than prefill in DeepSeek (Liu et al., 2024a)) creates additional offloading opportunities by allowing experts to be compacted onto fewer devices.
Adaptation to other MoE architectures.. FineMoE can be easily integrated with different MoE architectures. For shared experts, we treat them as always-hit during expert prediction. One of our evaluated models, Qwen1.5-MoE, includes shared experts that are used by all tokens. For multi-gating MoE, we can extend the expert map search by recording each gate’s probability distribution and flattening the outputs into a single vector for efficient similarity computation. This enables unified handling across diverse routing schemes.
8. Related Work
In this section, we provide a brief overview of recent studies and related works on MoE serving.
@@ 8. Related Work
Lossless MoE serving. Recent studies on lossless MoE serving have been widely proposed. DeepSpeed Inference (Aminabadi et al., 2022) offloads layer-wise parameters without considering expert awareness and does not provide expert prefetching or caching capabilities. Mixtral-Offloading (Eliseev and Mazur, 2023) employs LRU expert caching and introduces speculative prediction to enable expert prefetching. MoE-Infinity (Xue et al., 2024) proposes the request-level expert activation matrix to guide offloading in coarse granularity. SwapMoE (Rui Kong, Yuanchun Li, Qingtian Feng, Weijun Wang, Xiaozhou Ye, Ye Ouyang, Linghe Kong, Yunxin Liu, 2023) maintains a set of critical experts in GPU memory and adjusts them based on workload changes to minimize offloading overhead. ProMoE (Song et al., 2024) trains predictors per MoE layer to achieve high speculative prediction accuracy and low inference latency. Lina (Li et al., 2023) exports unpopular experts to host memory while focusing on MoE training. Liu et al. (2025) partitions and serves MoE models on serverless computing. Fiddler (Kamahori et al., 2025) serves MoE inference on CPU and GPU collaboratively. MoEShard (Balmau et al., 2025) shards experts to achieve balanced expert loads. Unlike existing coarse-grained offloading solutions, FineMoE tracks fine-grained expert patterns from both trajectory and semantic aspects and outperforms SOTA baselines.
Lossy MoE serving. Expert pruning (Dmitry Lepikhin, HyoukJoong Lee, Yuanzhong Xu, Dehao Chen, Orhan Firat, Yanping Huang, Maxim Krikun, Noam Shazeer, Zhifeng Chen, 2021) reduces memory usage by removing under-utilized experts. Expert compression (Pingzhi Li, Zhenyu Zhang, Prateek Yadav, Yi-Lin Sung, Yu Cheng, Mohit Bansal, Tianlong Chen, 2024; Tang et al., 2024; Zhou et al., 2025; Wu et al., 2025; He et al., 2024) compresses less-popular experts to reduce models’ memory footprint. Expert load rerouting (Gupta et al., 2024; Zhang et al., 2025; He et al., 2025) balances tokens to under-loaded experts instead of following the outputs of gate networks. Specifically, Hobbit (Tang et al., 2024) uses low precision to serve less-critical experts. He et al. (2025) drops and reroutes tokens from overloaded experts to others to reduce the straggler effect. Samoyeds (Wu et al., 2025) serves MoE models with sparsity computing. Lynx (Gupta et al., 2024) selects experts based on batch-level expert importance instead of gate network outputs. DAOP (Zhang et al., 2025) performs computations with predicted experts directly and cannot guarantee full generation quality. FLoE (Zhou et al., 2025) compresses experts on-the-fly during inference. However, lossy serving impacts the model quality and is orthogonal to FineMoE.
@@
MoE refactorization. Some works propose to redesign and refactor the current MoE architecture, such as decoupling gate networks from inference process (Hwang et al., 2024; Du et al., 2024) or building activation-efficient MoE models (Cai et al., 2024; Jie et al., 2025). This line of work requires model training or fine-tuning before serving. Pre-gated MoE (Hwang et al., 2024) trains pre-gate functions to eliminate the sequential dependencies between expert selection and execution. SiDA (Du et al., 2024) proposes a sparsity-inspired data-aware inference system that decouples the expert routing from inference. READ-ME (Cai et al., 2024) refactors pre-trained dense LLMs into specialized MoE models. MoLE (Jie et al., 2025) replaces inputs of all MoE layers with embedding tokens to avoid sparse expert activation. In contrast, FineMoE requires zero training to serve MoE models.
9. Conclusion